OpenStudy (anonymous):

Solve the equation. sqrt 3x-5=4

OpenStudy (anonymous):

square both sides 3x- 5 = 4^2 3x- 5 = 16 3x = 16 +5 3x = 21 x = 21/3 x = 7

OpenStudy (anonymous):

@Lindsayb12 you keep forgetting the sqrt

OpenStudy (anonymous):

o okay